# A Copula-Based Variational Autoencoder for Uncertainty Quantification in Inverse Problems: Application to Damage Identification in an Offshore Wind Turbine

### v1

@@ Abstract
Structural Health Monitoring of Floating Offshore Wind Turbines (FOWTs) is critical for ensuring operational safety and efficiency. However, identifying damage in components like mooring systems from limited sensor data poses a challenging inverse problem, often characterized by multimodal solutions where various damage states could explain the observed response. To overcome it, we propose a Variational Autoencoder (VAE) architecture, where the encoder approximates the inverse operator, while the decoder approximates the forward. The posterior distribution of the latent space variables is probabilistically modeled, describing the uncertainties in the estimates. This work tackles the limitations of conventional Gaussian Mixtures used within VAEs, which can be either too restrictive or computationally prohibitive for high-dimensional spaces. We propose a novel Copula-based VAE architecture that decouples the marginal distribution of the variables from their dependence structure, offering a flexible method for representing complex, correlated posterior distributions. We provide a comprehensive comparison of three different approaches for approximating the posterior distribution: a Gaussian Mixture with a diagonal covariance matrix, a Gaussian Mixture with a full covariance matrix, and a Gaussian Copula model. Our analysis, conducted on a high-fidelity synthetic dataset, demonstrates that the Gaussian Copula VAE offers a promising and tractable solution in high-dimensional spaces. Although the present work remains in the two-dimensional space, the results suggest efficient scalability to higher dimensions. It achieves slightly superior performance with significantly fewer parameters than the Gaussian Mixture alternatives, whose parametrization grows prohibitively with the dimensionality of the latent space. The results underscore the potential of Copula-based VAEs as a powerful tool for uncertainty-aware damage identification in FOWT mooring systems, generalizable to other structural systems.
@@ 1 Introduction
Offshore wind energy solutions have become increasingly appealing for the green transition, owing to the energy production opportunities offered by the strong winds in these deep-water sites [1]. Proof of this interest is the 117 GW of new wind energy installed worldwide in 2024 [2]. However, these advantageous conditions come along with drawbacks, such as harsh operating conditions, the reduced accessibility [3], and the need to deploy complex mooring and floating systems as traditional fixed-bottom devices become unfeasible[4]. These challenging characteristics result in expensive maintenance plans that expose workers to hazardous environments during the inspections and repair interventions [5].
@@
The mooring system is one of the most critical components in Floating Offshore Wind Turbines (FOWTs). It is responsible for ensuring adequate anchorage to the seabed and limiting the platform’s movements in the presence of deep currents, strong wind speeds, and waves. The harsh, corrosive seawater environment in which these assets operate makes the mooring system highly vulnerable. The adherence of algae to the mooring lines (biofouling) [6] and the displacement or trawling of the anchor (anchoring) [7] are two frequently occurring damage types that need attention. If these phenomena remain undetected over a prolonged period, the excessive, uncontrolled motion of the platform may compromise the integrity of the FOWT and reduce its power production capability.
@@
In this context, the role of Structural Health Monitoring (SHM) for continuously and remotely assessing the state of FOWTs gains relevance. Framed as an inverse problem, the goal of SHM is to infer the condition of the mooring system from a limited set of sensors that measure its response. The instrumentation devices range from low-cost, low-maintenance, and easily deployable sensors (e.g., accelerometers on the tower or inclinometers on the platform) to more complex and expensive ones (e.g., direct tension mooring line measurement)[8]. Over the last decade, significant efforts have been made to leverage these data, ranging from statistical pattern recognition tools [9] to Artificial Intelligence (AI) [10, 11].
The most extensively employed AI-based tool are Deep Neural Networks (DNNs) [12], which have been increasingly settled in the field of FOWT assessment in the past five years. For example, work [13] employs a fully-connected DNN to detect mooring line damage in a tension leg platform using statistical features of the floater’s six Degrees Of Freedom (DOFs). They implement a classification architecture to identify nine potential damage scenarios, including the healthy state. In [14], which served as the starting point for this paper, the authors implement a DNN to identify two damage types (biofouling and anchoring) in a mooring line of a FOWT using statistics from the six platform’s rigid body DOFs. Standard fully connected DNNs have given way to Convolutional and Recurrent NNs [15, 16] for efficiently handling multichannel time series. For example, in [17], Sharma et al. propose a novel Convolutional Neural Network (CNN) architecture with an embedded autoregressive model to compress and enhance the expressiveness of the data. Work [18] fuses time-frequency images from rotational motion DOFs to feed a Recurrent Neural Network (RNN) for anomaly detection in the mooring lines of an offshore semisubmersible platform. Recently, Long Short-Term Memory (LSTM) architectures have been employed. Work [19] combines an LSTM layer with Support Vector Machines to exploit time series in the task of damage classification. Most of these works are validated in benchmark experiments or numerical studies, often accounting for the effect of Gaussian noise via percentage increase mechanisms [13, 20, 18]. Work [21] extends the analysis to impulse noise and signal-dependent noise when identifying damage in a floating platform.
However, in real applications, the uncertainty sources are multiple and diverse, including environmental and operational variability, measurement error, and modeling error, each of a different nature and behavior. Above all, there is one additional challenge preventing the well-posedness of this inverse problem: multimodality. Multimodality refers to the existence of multiple plausible solutions that arise from slight perturbations or errors in the measured data. This situation exacerbates as the instrumentation system is reduced. In FOWTS, due to practicality issues, scarce and low-accuracy sensors are often used for long-term monitoring. Under such circumstances, deterministic data-driven strategies may drastically fail in their diagnostics, often yielding physically meaningless outcomes.
The core pitfall of standard architectures to reliably solve inverse problems is the loss function, which may exhibit multiple modes and spurious local minima. This situation makes the solution strongly dependent upon the initialization assumptions [22]. In work [23], the authors analyze the limitations of expressing the loss in the ill-posed space of the inversion outcomes. They explore an encoder-decoder architecture to build the loss in the measurement space, thus ensuring the physical consistency of the estimates. But to fully describe the solution space of the inverse problem, probabilistic approaches are needed. Variational Autoencoders (VAEs) bring a powerful solution to tackle this problem [24]. Compared to standard AEs, VAEs embed a distributional model to statistically describe the latent space posterior distribution given some prior information [25]. VAEs have been recently adopted to solve inverse problems, given their capability to represent the uncertainty in the solution space. Work [26] introduces the use of VAEs for uncertainty quantification in scientific inverse problems. In work [27], the authors apply a Convolutional VAE to account for uncertainty in the long-term fatigue estimations of a wind turbine blade. The recent work [28] extends the loss function in [23] to account for the posterior distribution discrepancy of the estimated resistivity properties in a magnetotelluric inverse problem.
In our previous work [29], we followed the methodology in [28] to tackle uncertainty in the inverse problem of damage identification originally introduced in [14]. We employed Gaussian Mixtures (GMs) to represent the posterior distribution of a two-dimensional latent space describing the mooring line damage condition: biofouling and anchoring levels. Despite the successful results obtained at the validation stage, the authors identified a critical limitation that required further analysis: the strongly restrictive assumption of diagonal covariance matrices in the GM components. Such an assumption restricted the number of parameters to be estimated at the cost of extremely simple distributional shapes. This work contributes to this line by deeply exploring and comparing more suitable distributional models in the VAE architecture. We first extend the GM approach proposed in [29] to consider full-covariance matrices, endowing the distributional model with more flexibility to capture complex posterior distributions. One key shortcoming of this approach is the prohibitive increase in the number of parameters as the dimensionality of the latent space increases[25].
@@
Herein, we propose a Copula-based distributional model [30] to tackle the scalability limitations of full-covariance GMs. Copulas are functions that join the multivariate distribution (describing the interrelation among variables) to the independent one-dimensional marginal distributions in an efficient manner [31, 32]. Copulas were originally introduced by Sklar [33], who defined in his theorem the properties of Copula functions. They can adopt various forms, mainly divided into (a) parametric (e.g., Gaussian, Archimedean), or (b) non-parametric (e.g., Clayton, Gumbel) [34, 35]. Copulas have been widely employed in the field of finance and risk management [36, 37, 38]. In image reconstruction, work [39] employs a Gaussian Copula to model the local dependency in mixed categorical and continuous data. In language modeling, work [40] proposed a Gaussian Copula-based VAE architecture to efficiently capture the dependency relationships over latent variables. They applied and validated the methodology using three benchmark vocabulary datasets. Recently, work [41] used the Copula model to maintain the dependency among latent variables in a collaborative filtering VAE; and [42] proposed a Contrastive Copula VAE to eliminate coupled representations in a self-supervised classifier.
@@
However, Copulas remain unexplored in the field of inverse problems, particularly for condition assessment of FOWTs. Thus, another contribution of this work is the use of a Gaussian Copula to describe the posterior distribution of the mooring line’s damage features given measurements from its response. We design a Gaussian Copula-based VAE architecture that includes a differentiable sampling strategy to feed the decoder with randomly drawn samples. This work provides a proof of concept for the two-dimensional case study presented in [29], comparing its performance and scalability with those of the diagonal covariance and full-covariance GM approaches. Although the present work remains in the two-dimensional space, the results promise a tractable and efficient scalability to higher dimensions. The analysis reveals the benefits of using Copulas to describe high-dimensional interrelated spaces, a common scenario in the field of advanced damage identification (location and quantification) from scarce response measurements.
@@
We employ statistics from short-term motion time series generated in a synthetic environment using OpenFAST software [43]. The lack of labeled experimental data from an operating FOWT imposes this limitation on our work. Nonetheless, the simulator is reliable and produces high-fidelity, realistic data. Another limitation is that we consider static damage, neglecting the time-domain evolution of the system’s condition, required for fatigue analysis and Remaining Useful Life estimates, which is considered future work. Finally, our method robustly expresses and represents the multi-source uncertainty in the inverse outcome estimates, but it is unable to disentangle the different sources (measurement error, modeling error, scarce sensing, etc.).
@@
The rest of the paper is organized as follows: Section 2 (a) describes the physical problem to solve, including the FOWT dynamics and the data generation process; (b) formulates the VAE architecture and associated loss function; and (c) expresses the different posterior distributions to be explored (i.e., the diagonal and full covariance Gaussian Mixtures, and the Gaussian Copula), including the sampling strategy. Section 3 reports the specifications to build the VAE architecture, including the implementation and training details. Section 4 presents the main results in the form of a comparative analysis. Finally, Section 5 summarizes the conclusions and future work.
@@ 2.1 Governing Dynamics of FOWTs
$(M+A_{\infty})\ddot{q}(t)+Kq(t)=\sum F_{\text{ext}}(t,\omega),$ (1)
where $q(t)$ contains the six rigid-body degrees of freedom (DOFs): surge, sway, heave, roll, pitch, and yaw, and $\ddot{q}(t)$ is the corresponding vector of accelerations. Matrices $M$ and $K\in\mathbb{R}^{n_{\text{dof}}\times{n_{\text{dof}}}}$ encapsulate the structural mass and stiffness of the FOWT system, respectively, for the six DOFs ($n_{\text{dof}}=6$). Matrix $A_{\infty}\in\mathbb{R}^{n_{\text{dof}}\times n_{\text{dof}}}$ represents the added mass at infinite frequency, which describes the inertial effect of the irrotational ideal fluid that is accelerated when the platform experiments unsteady motion. The right-hand term comprises all the external forces that affect the system, which depend upon time $t$ and the wave frequency $\omega$ The external forces can be decomposed into various components, including hydrodynamic, aerodynamic, and radiation damping, among others.
@@
$(M_{S}+A_{\infty})\ddot{q}(t)+\int_{0}^{t}R_{S}(t-\tau)\dot{q}(\tau)d\tau+K_{S}q(t)=F_{\text{wave}}(t,\mathbf{w})+F_{\text{wind}}(t,\mathbf{w})+F_{\text{visc}}(t,q,\dot{q})+F_{\text{moor}}(t,q,\mathbf{w},S_{\text{moor}}),$ (2)
where $F_{\text{wave}}(t,\mathbf{w})$ denotes the wave-induced forces, $F_{\text{wind}}(t)$ denotes the aerodynamic forces, $F_{\text{visc}}$ denotes the viscous hydrodynamic forces and moments owing to fluid viscosity, and $F_{\text{moor}}(t,q,\mathbf{w},S_{\text{moor}})$ denotes the mooring system forces. Subscript $S$ denotes the system properties, and $\mathbf{w}$ denotes the environmental loading conditions. The retardation functions $R_{S}(t)$ are related to the frequency-dependent radiation damping coefficients $B(\omega)$ via Fourier transform:
@@
The hydrodynamic coefficients $A(\omega)$ and $B(\omega)$ are typically pre-calculated using boundary element method (BEM) codes (e.g., WAMIT, AQWA). Equation 2 is the basis used by FOWT simulation tools such as OpenFAST [48].
The solution $q(t)$ to the governing motion equation contains the time-domain response of the FOWT. In real practice, only some of the DOFs are measured. The most frequently measured signals often correspond to the rotations, as clinometers provide an accurate and economic technology. Changes in mooring line integrity directly influence the terms representing mooring force $F_{moor}(t,q,S_{moor})$ in Eq. 2. This alteration of mooring system properties $S_{moor}$ manifests in the system’s response. According to [29], in the following, we describe the selected relevant features and how mooring damage affects them. We first obtain the mean displacement as:

### v2

@@ Abstract
Structural health monitoring of floating offshore wind Turbines (FOWTs) is critical for ensuring operational safety and efficiency. However, identifying damage in components like mooring systems from limited sensor data poses a challenging inverse problem, often characterized by multiple solutions where various damage states could explain the observed response. To overcome this, we propose a variational autoencoder (VAE) architecture, where the encoder approximates the inverse operator that maps the observed response to the system’s condition, while the decoder approximates the forward operator that maps the system’s condition and measured excitation to its response. Conventional Gaussian mixture models can be restrictive and are often computationally prohibitive when used within VAE. This work addresses these limitations by proposing a novel copula-based VAE architecture that decouples the marginal distribution of variables from their dependence structure, providing a flexible method for representing complex, correlated posterior distributions. In our method, the observed response corresponds to statistical features derived from short-term rotation motion signals of the FOWT platform. The damage condition is described by the severity level of two damage classes frequently found in the mooring system (anchoring and biofouling). We provide a comprehensive comparison of the copula against standard Gaussian and Gaussian mixture approaches, considering both diagonal and full covariance matrices, to demonstrate the benefits of copulas. Our analysis, conducted on a high-fidelity synthetic dataset, demonstrates that the Gaussian Copula VAE offers a promising and tractable solution in high-dimensional spaces. Although 2D, the number of copula parameters grows much more slowly with the dimension than the other methods, and therefore it shows promise for higher-D scalability. In the test experiments, the copula achieves superior performance with significantly fewer parameters than the Gaussian Mixture alternatives, whose parametrization grows prohibitively with the dimensionality of the latent space. The results highlight the potential of copula-based VAEs as a powerful tool for uncertainty-aware damage identification in FOWT mooring systems.
@@ 1 Introduction
Continuously assessing the health condition of the mooring system in floating offshore wind turbines (FOWTs) is crucial to ensure their durability and enhance their contribution to the power network. This paper proposes an efficient copula-based variational autoencoder (VAE) to identify damage in the mooring system of an FOWT from limited motion responses measured through monitoring. By designing a copula-based posterior distribution, we efficiently express the multiplicity of solutions to the inverse problem. Once trained, the VAE receives new motion observations and delivers a map of the feasible solutions that most likely produced them, describing the uncertainty of the diagnostic.
@@
Offshore wind energy solutions offer outstanding energy production opportunities thanks to the strong winds that exist in the deep-water sites where they are installed [1]. As a consequence, they have become increasingly attractive for the European green transition [2]. Indeed, 117 GW of new wind power capacity was installed globally in 2024 only [3]. However, these appealing characteristics also bring drawbacks, such as harsh operating conditions, reduced accessibility [4], and complexity of the required mooring and floating systems [5]. These challenging characteristics result in expensive and potentially unsafe maintenance operations, which may expose workers to hazardous environments during inspection and repair intervention [6].
@@
Mooring systems are one of the most critical components in floating offshore wind turbines (FOWTs). They are responsible for ensuring adequate anchorage to the seabed and limiting the platform’s movements in the presence of deep currents, strong wind speeds, and waves. The harsh, corrosive seawater environment in which these assets operate makes the mooring system highly vulnerable. In fact, multiple types of damage frequently occur. Among these, the most frequent are biofouling, caused by the adherence of algae or other organisms [7], and anchoring, i.e., the trawling of the mooring system anchor [8]. If these phenomena remain undetected over a prolonged period, they may cause the platform to move excessively, thereby compromising the FOWT’s integrity and its power production capability.
@@
In this context, structural health monitoring (SHM) is vital for continuously and remotely assessing the current state of FOWTs’ components. Its goal is to infer the condition of the mooring system from a limited set of sensors that measure its response, and thus can be posed as an inverse problem. A wide range of sensors is available, from low-cost, low-maintenance, and easily deployable sensors (e.g., accelerometers on the tower or inclinometers on the platform) to more complex and expensive ones (e.g., direct tension mooring line measurement) [9]. Over the years, significant efforts have been made to exploit these data, ranging from statistical pattern recognition tools [10] to artificial intelligence (AI) [11, 12].
Deep neural networks (DNNs) [13] have become extensively employed in the field of FOWT assessment, from standard fully-connected to convolutional, recurrent, or long-short-term-memory (LSTM) approaches [14, 15, 16, 17, 18, 19]. However, in real applications, there exist multiple uncertainty sources, including environmental and operational variability, measurement and modeling error, or data limitations (reduced instrumentation system with a limited sensitivity to the target damage), each with a different nature and behavior. Under such circumstances, the solution of the inverse problem becomes non-unique, meaning that slight perturbations may result in extremely different outcomes. When training deterministic DNNs to solve such ill-posed inverse problems, the non-convexity of the loss functions typically employed makes the estimated solution strongly dependent upon the initialization assumptions [20]. This phenomenon seriously compromises the reliability of the method and, in consequence, the associated decision-making in SHM applications [21, 22]. The authors of [23] analyze the limitations of expressing the loss in the space of the inversion outcomes. They explore an autoencoder (AE) architecture that builds the loss in the measurement space, thus ensuring the physical consistency of the inverse estimates.
Despite the benefits of AE architectures, probabilistic approaches are needed to fully describe the solution space of the inverse problem [24]. Variational autoencoders (VAEs) bring a powerful solution to tackle uncertainty in the inference phase [25]. Compared to standard AEs, VAEs embed a distributional model to statistically describe the latent space posterior distribution given some prior information [26]. They rely on Bayesian inference theory to account for the discrepancy between the estimated and the true posterior distributions, assuming a known error distribution [25].
VAEs have been extensively employed as generative models with dimensionality reduction and clustering purposes [27, 28, 29, 30, 31]. The works by Goh et al. [32] and Almaeen et al. [33] introduced the use of VAEs for uncertainty quantification in scientific inverse problems. Since then, various works have exploited the potential of VAEs in different inversion applications [34, 35, 36, 37, 38, 39]. In the context of SHM, VAEs have been mostly applied for advanced damage-sensitive feature extraction to detect structural changes from measured response data [40, 41, 42, 43]. Simpson et al. [44] and Bacsa et al. [45] employ LSTM VAEs to infer changes in the material properties of a multi-degree-of-freedom (MDOF) structure. In [46], the authors employ a VAE to classify different structural damage types by applying a clustering to the latent space. The authors of [47] apply a Convolutional VAE to account for uncertainty in the long-term fatigue estimations of a wind turbine blade. However, all these works assume a normal (Gaussian) posterior probability density function (PDF) with a diagonal covariance matrix, which imposes strong constraints to the posterior PDF shape and prevents a complete description of the uncertainty effect.
In order to improve the expressivity of VAEs, more complex parametric posterior PDFs have been explored, with particular interest in the use of mixtures [48]. In [49], the authors employ a Gaussian mixture to describe the latent space for image reconstruction. Liu et al. [50] apply a sequence of invertible Householder transformations to a Gaussian mixture to achieve arbitrarily complex posteriors. In [51], the authors extend the loss function from [23], proposing a diagonal covariance-Gaussian mixture latent posterior PDF to describe uncertainty in the estimated resistivity properties. The same approximate posterior PDF is used in [52] to identify a two-dimensional damage condition in the mooring system of an FOWT. Although employing a mixture significantly broadens the expressivity compared to single normal approaches, the strongly restrictive assumption of a diagonal covariance matrix demands a high number of components to describe complex distributional shapes. As a consequence, the number of parameters to be estimated increases, making this approach hardly scalable beyond two-dimensional latent spaces.
@@
An efficient approach to describe complex multivariate distributions is the copula [53]. Copulas are functions that join the multivariate distribution (describing the interrelation among variables) to the independent one-dimensional marginal distributions [54, 55]. Copulas have been widely employed in the field of finance and risk management [56, 57, 58]. In image reconstruction, work [59] employs a Gaussian copula to model the local dependency in mixed categorical and continuous data. In language modeling, Wang et al. [60] propose a Gaussian copula-based VAE architecture to efficiently capture the dependency relationships over latent variables. Zhong et al. [61] use the copula posterior PDF to preserve dependency among latent variables in a filtering VAE. Wu et al. [62] employ a contrastive copula VAE to isolate and eliminate coupled dependencies in a self-supervised classifier.
@@
Despite the increasing interest in copula-based models to enhance VAEs’ performance [63], this mechanism remains unexplored in the field of SHM. To the authors’ knowledge, this is the first work to embed a copula in a VAE for approximating the multivariate posterior PDF and inferring the damage condition of a structural system. We first extend the Gaussian mixture approach introduced in [52], considering full-covariance matrices to enhance expressivity with respect to the diagonal-covariance simplification. The key shortcoming of this approach is the prohibitive increase in the number of parameters as the dimensionality of the latent space increases [26]. To overcome this scalability limitation, we then propose a copula-based distributional model [53]. We design a Gaussian copula VAE architecture that includes a differentiable rejection sampling strategy to feed the decoder with randomly drawn samples within the desired latent domain. We implement the methodology as a proof of concept for the two-dimensional case study presented in [52], which aims to infer the presence of damage in the mooring system of an FOWT from motion measurements of its platform. We exhaustively compare the performance of the copula with that of the diagonal- and full-covariance Gaussian mixture approaches. Although the present work remains in the two-dimensional space, the results obtained with the copula promise a tractable and efficient scalability to higher dimensions. The analysis reveals the benefits of using copulas to describe high-dimensional interrelated spaces, a common scenario in the field of advanced damage identification (location and quantification) from scarce response measurements.
@@
We employ statistics from short-term motion time series generated in a synthetic environment using the OpenFAST open-source software library [64]. The lack of labeled experimental data from an operating FOWT imposes this limitation on our work. Nonetheless, the simulator is reliable and produces high-fidelity, realistic data owing to the extensive research and validation efforts using experimental data [65, 66, 67, 68, 69]. Another limitation is that we consider static damage, neglecting the time-domain evolution of the system’s condition, which is required for fatigue analysis and remaining useful life estimates, and is considered future work. Finally, our method robustly expresses and represents the multi-source uncertainty in the inverse outcome estimates, but it is unable to disentangle the different sources (measurement error, modeling error, sensing scarcity, etc.).
@@
The rest of the paper is organized as follows: Section 2 (a) describes the physical problem to solve, including the FOWT dynamics and the data generation process; (b) formulates the VAE architecture and associated loss function; and (c) expresses the different posterior distributions to be explored (i.e., the diagonal and full covariance Gaussian mixtures, and the Gaussian copula), including the sampling strategy. Section 3 reports the specifications to build the VAE architecture, including the implementation and training details. Section 4 presents the main results in the form of a comparative analysis. Finally, we draw our conclusions and describe opportunities for future work in Section 5.
@@ 2.1 Governing Dynamics of FOWTs
$(M+A_{\infty})\ddot{u}(t)+Ku(t)=\sum F_{\text{ext}}(t,\omega),$ (1)
where $u(t)$ contains the six rigid-body degrees of freedom (DOFs): surge, sway, heave, roll, pitch, and yaw, and $\ddot{u}(t)$ is the corresponding vector of accelerations. Matrices $M$ and $K\in\mathbb{R}^{n_{\text{dof}}\times{n_{\text{dof}}}}$ encapsulate the structural mass and stiffness of the FOWT system, respectively, for the six DOFs ($n_{\text{dof}}=6$). Matrix $A_{\infty}\in\mathbb{R}^{n_{\text{dof}}\times n_{\text{dof}}}$ represents the added mass at infinite frequency, which describes the inertial effect of the irrotational ideal fluid that is accelerated when the platform experiments unsteady motion. The right-hand term comprises all the external forces that affect the system, which depend upon time $t$ and the wave frequency $\omega$ The external forces can be decomposed into various components, including hydrodynamic, aerodynamic, and radiation damping, among others.
@@
$(M_{S}+A_{\infty})\ddot{u}(t)+\int_{0}^{t}R_{S}(t-\tau)\dot{u}(\tau)d\tau+K_{S}u(t)=F_{\text{wave}}(t,\mathbf{w})+F_{\text{wind}}(t,\mathbf{w})+F_{\text{visc}}(t,u,\dot{u})+F_{\text{moor}}(t,u,\mathbf{w},S_{\text{moor}}),$ (2)
where $F_{\text{wave}}(t,\mathbf{w})$ denotes the wave-induced forces, $F_{\text{wind}}(t)$ denotes the aerodynamic forces, $F_{\text{visc}}$ denotes the viscous hydrodynamic forces and moments owing to fluid viscosity, and $F_{\text{moor}}(t,u,\mathbf{w},S_{\text{moor}})$ denotes the mooring system forces. Subscript $S$ denotes the system properties, and $\mathbf{w}$ denotes the environmental loading conditions. The retardation functions $R_{S}(t)$ are related to the frequency-dependent radiation damping coefficients $B(\omega)$ via Fourier transform:
@@
The hydrodynamic coefficients $A(\omega)$ and $B(\omega)$ are typically pre-calculated using boundary element method (BEM) codes (e.g., WAMIT, AQWA). Equation (2) is the basis used by FOWT simulation tools such as OpenFAST [74].
The solution $q(t)$ to the governing motion equation contains the time-domain response of the FOWT. In real practice, only some of the DOFs are measured. The most frequently measured signals often correspond to the rotations, as clinometers provide an accurate and economic technology. Changes in mooring line integrity directly influence the terms representing mooring force $F_{moor}(t,u,S_{moor})$ in eq. (2). This alteration of mooring system properties $S_{moor}$ manifests in the system’s response. According to [52], in the following, we describe the selected relevant features and how mooring damage affects them. We first obtain the mean displacement as: